# 1]FAIR at Meta Dualformer: Controllable Fast and Slow Thinking by Learning with Randomized Reasoning Traces

### v2

@@ Abstract
In human cognition theory, human thinking is governed by two systems: the fast and intuitive System 1 and the slower but more deliberative System 2. Recent studies have shown that incorporating System 2 process into Transformers including large language models (LLMs), significantly enhances their reasoning capabilities. Nevertheless, models that purely resemble System 2 thinking require substantially higher computational costs and are much slower to respond. To address this challenge, we present Dualformer, a single Transformer model that seamlessly integrates both the fast and slow reasoning modes. Dualformer is obtained by training on data with randomized reasoning traces, where different parts of the traces are dropped during training. The dropping strategies are specifically tailored according to the trace structure, analogous to analyzing our thinking process and creating shortcuts with patterns. At inference time, our model can be configured to output only the solutions (fast mode) or both the reasoning chain and the final solution (slow mode), or automatically decide which mode to engage (auto mode). In all cases, Dualformer outperforms the corresponding baseline models in both performance and computational efficiency: (1) in slow mode, Dualformer optimally solves unseen maze navigation tasks of the time, surpassing the Searchformer (trained on data with complete reasoning traces) baseline performance of 93.3%, while only using fewer reasoning steps; (2) in fast mode, Dualformer completes those tasks with an optimal rate, significantly outperforming the Solution-Only model (trained on solution-only data), which has an optimal rate of only 30%; (3) when operating in auto mode, Dualformer achieves an optimal rate of 96.6% while utilizing fewer reasoning steps compared to Searchformer. For math problems, our techniques have also achieved improved performance with LLM fine-tuning, showing its generalization beyond task-specific models.
@@ 1 Introduction
In psychology, the dual process theory Wason and Evans (1974) explains that thought can emerge through two distinct processes. Typically, one process is implicit (automatic and unconscious) and the other process is explicit (controlled and conscious).
The famous book Thinking, Fast and Slow Kahneman (2017) explores the concept of dual process theory extensively. It describes two different systems of thinking. System 1, corresponding to the implicit process, is fast, intuitive, and emotional. System 2, corresponding to the explicit process, is slower, more deliberative, and more logical. Kahneman applies this theory to a wide range of cognitive and behavioral activities to illustrate how these two systems influence our decision making and reasoning. As outlined in the book, System 1 can lead to biases and is prone to systematic errors due to the lack of voluntary control. As a counterbalance, System 2 is more suitable for tasks that require careful analysis, consideration, and planning. 00footnotetext: The code is available at https://github.com/facebookresearch/dualformer.
@@
Transformers Vaswani et al. (2017), the sequence modeling tool that serves as the cornerstone of foundation models in various domains including Large Language models (LLMs) Dosovitskiy (2020); Baevski et al. (2020); Radford et al. (2021); Touvron et al. (2021); Hsu et al. (2021); Touvron et al. (2023); Dubey et al. (2024), have been widely used in many works to approach reasoning and planning problems, see e.g., Zhou et al. (2022); Kojima et al. (2022); Pallagani et al. (2022); Valmeekam et al. (2023a); Chen et al. (2024); Gundawar et al. (2024); Wang and Zhou (2024). Interestingly, the dual process theory generalizes to artificial intelligence, where we can categorize the reasoning modes of Transformer into fast and slow. In fast mode, a Transformer will directly output a final solution without any reasoning steps, whereas the intermediate steps of thinking, such as a search trace for finding a short path, will be generated along with the plan in slow mode. The two inference modes share similar pros and cons with the two thinking systems inherent in us. As discussed in previous works Wei et al. (2022); Valmeekam et al. (2023b); Lehnert et al. (2024); Gandhi et al. (2024); Saha et al. (2024), models operating in fast mode have a lower computational cost and allow for a quicker response, yet they fall short in accuracy and optimality compared with slow mode models.
This raises the question:
@@
Can we integrate both the fast and the slow modes into Transformer-based reasoning agents, similar to how humans possess two distinct thinking systems, and let them complement each other?
@@
Multiple approaches have been proposed. A popular paradigm is to start with a pure System 2, and then finetune to make it more efficient like System 1, e.g., to distill System 2 output into System 1 Yu et al. (2024); Deng et al. (2024), or to improve the reasoning efficiency of existing system 2 by distillation Wang et al. (2023); Shridhar et al. (2023) or bootstrapping from symbolic systems (e.g., Searchformer Lehnert et al. (2024), Stream of Search Gandhi et al. (2024)). However, in these cases, further fine-tuning is needed, which is computationally expensive, and it is non-trivial to adapt the resulting system to be more like System 1 or System 2 on the fly. To address this problem, Saha et al. (2024) design an explicit meta-controller to switch between two different systems.
@@
We first apply our framework to train an encoder-decoder Transformer model to solve pathfinding problems, where the trace is generated by the A* search algorithm. We consider two domains: the Maze navigation and the Sokoban game as in Lehnert et al. (2024), where we use the same tokenization scheme. Interestingly, we have found that these problems are challenging for state-of-the-art LLMs like o1-preview and o1-mini, where the output path often breaks into the walls (See Section 13 for an example).
In each reasoning mode, Dualformer outperforms the established baselines, achieving stronger results in both the solved rate and optimal rate. Moreover, Dualformer significantly enhances the diversity of the generated plans by identifying a greater variety of unique paths that reach the goal. Notably, Dualformer is also efficient even when working in slow mode, generating much shorter reasoning traces than the baseline model. Next, we apply our framework to finetune LLMs for answering math questions. Following Yu et al. (2023), the training examples are taken from the MATH dataset Hendrycks et al. (2021) where answers are rewritten by a Llama-3.1-70B-Instruct model to include detailed intermediate steps. Likewise, the obtained LLMs demonstrate enhanced efficacy and efficiency.
@@ 2 Related Work
Tremendous efforts have been made to enhance the capability of Transformer-based models to plan and reason over a long horizon. Two main types of approaches have been developed. The first type leverages existing LLMs. For instance, researchers taught LLMs to call external existing symbolic solvers, such as those found in Ahn et al. (2022); Besta et al. (2024); Sel et al. (2023); He-Yueya et al. (2023); Liu et al. (2023); Silver et al. (2024). Pallagani et al. (2022, 2024) investigate the fine-tuning of LLMs to use a symbolic solver, Schick et al. ; Hao et al. (2024) fine-tune LLMs to use external tools, and Hao et al. propose to fine-tune LLMs to do reasoning and planning within a world model. Among them, several works try to integrate System-1 and System-2 thinking into LLM reasoning. Weston and Sukhbaatar (2023) proposed the System 2 Attention (S2A), a more deliberate attention mechanism aimed at improving the reasoning capabilities of large language models by reducing their susceptibility to spurious correlations in the context. Yu et al. (2024) distill System 2 output into a System 1 model, which aims to compile higher-quality outputs from System 2 techniques back into LLM generations without intermediate reasoning token sequences.
@@
Large language models (LLMs) have been utilized for synthetic data generation in various domains. For instance, Wei et al. (2021); Longpre et al. (2023); Taori et al. (2023) introduced a synthetic instruction dataset by sampling from diverse templates containing natural language instructions that outline specific tasks. This approach has also been applied to the visual domain Liu et al. (2024b, a); Zhu et al. (2023); Brooks et al. (2023); Peng et al. (2023). To improve the performance of LLM to answer math questions, Yu et al. (2023) developed a method to rewrite, verify, and augment the original MATH dataset Hendrycks et al. (2021) using specialized prompts. Similar methodologies have been further explored in other studies, including those by Yuan et al. (2023); Luo et al. (2023); Lee et al. (2023); Yue et al. (2023); Tong et al. (2025).
3 Preliminaries
Our work builds upon the work of Lehnert et al. (2024). To perform planning, we train a Transformer to model a sequence of tokens that sequentially represents the planning task, the computation of A* algorithm, and the optimal solution derived from the A* search. The tokenization method is illustrated in Figure 3.1. As a toy example, we consider a navigation task in a Maze where the goal is to find one shortest path from the start cell to goal cell, without hitting a wall cell. The A* algorithm has successfully determined an optimal plan. We use a sequence of tokens to express both the task and the Maze structure, which also serves as the prompt for Dualformer. The solution is depicted by the plan token sequence that describes the path using coordinates. The A* algorithm generates a search trace sequence that records the search dynamics performed, displayed in Figure 4.1. Recall that the A* algorithm is a pathfinding algorithm on a weighted graph. The create clause adds the node (represented by the subsequent coordinates) into the search frontier, and the close clause adds the node to the closed set. Each clause, either create or close, is followed by the tokens x, y, c0, and c1, representing the node’s coordinates, cost-since-start value, and heuristic values, respectively. For details of A* and the tokenization approach, we refer the readers to Russell and Norvig (2021) and Lehnert et al. (2024), respectively.
@@ 4 Structured Trace Dropping and Randomized Training
Searchformer proposed by Lehnert et al. (2024) has proven efficacy in addressing a variety of complex decision-making tasks. However, it still suffers from two important limitations.
Firstly, the model only operates in slow mode and outputs lengthy reasoning chains, which significantly increases the inference time. While this can be reduced by bootstrapping Lehnert et al. (2024), an iterative refining technique that consists of cycles of rollouts followed by fine-tuning, such a procedure incurs significant extra demand on computational resources.
Secondly, Searchformer struggles to generate diverse solutions, where identical rollouts are frequently sampled111We note that the lack of diversity is not due to the limitation of training data. For all our experiments, the search traces in the training datasets are generated by non-deterministic A* algorithm that randomly breaks cost ties.. For example, across 1000 maze problems we have tested, Searchformer’s reasoning chain contains more than 1500 tokens on average, and can only find unique feasible paths out of responses (see Section 5.1.2).
To address these challenges, we propose a training framework that utilizes randomized reasoning traces. Our approach is inspired by two lines of work. First, we have noticed that even though Searchformer is trained on complete A* search traces, it generates shorter traces that are sketching the search process. Second, research has shown that humans often rely on shortcuts and patterns when making decisions, a concept known as System 1 thinking Kahneman (2017). These observations, combined with the success of dropout technique Hinton (2012); Srivastava et al. (2014) that randomly drop units from the neural network during training, motivated us to explore the use of randomized reasoning traces in our framework, and we aim to simplify the A* search trace by exploiting its structured elements and selectively dropping certain parts for each training example. The details of our approach are described below.
@@
the Level 1 strategy eliminates all the close clauses from a search trace.
The Level 2 strategy goes further by additionally dropping all the cost tokens.
The Level 3 strategy is even more aggressive. It further randomly drops of the create clauses.
Ultimately, the Level 4 strategy drops the entire trace.
@@
Curious readers might already be wondering that whether our training framework resembles the token masking techniques used by famous LLMs including BERT Devlin (2018); Liu (2019); Song (2019); Gauthier and Levy (2019); Sinha et al. (2021); Kitouni et al. (2024). However, there are significant differences that distinguish our approach from those masking techniques. First, standard masking techniques usually mask the input tokens of a sequence uniformly in random. In contrast, our dropping strategies only apply to the search trace. Second, while masked LLMs generally employ bidirectional attention layers and predict the masked tokens, Dualformer uses causal attention layers, and our training objective solely focuses on the next token prediction with the overall goal of improving its reasoning and planning capability. Computationally, our training procedure is also more efficient. Dropping tokens shortens the input sequence and saves computation. For example, it takes 30 hours to train Dualformer for the maze task on 8 Tesla V100 32GB GPUs, while it requires 36 hours if we use the full reasoning trace. We defer the training details to Section 5 and Appendix 7.
@@ 4.1 Controllable Generation
One appealing property of Dualformer is that it can be easily prompted to operate in either fast or slow generation mode at inference time. The control mechanism is extremely simple: we append bos and a control token to the standard prompt (which includes environment and task description), where the control token is either plan or create. If we use plan, Dualformer will operate in fast mode and directly output the plan, bypassing the reasoning steps. On the other hand, if we inject create after bos, Dualformer will work in slow mode and generate both the reasoning trace and the final plan. See Appendix 8 for concrete examples. If we only use the standard prompt, Dualformer will mimic the dual process of human decision making—depending on the situation, it generates either types of responses, which correspond to System 1 and System 2 reasoning.
@@ 5 Experiments
In slow model, does Dualformer lead to faster reasoning, i.e., output a shorter trace?
@@
We answer questions 1 and 2 in Section 5.1, where we train Transformers to solve Maze navigation tasks and the closely related Sokoban games, similar to SearchFormer Lehnert et al. (2024). To answer questions 3, we finetune LLama-3.1-8B and Mistral-7B models to solve math problems in Section 5.2.
@@ 5.1 Navigation Tasks: Maze & Sokoban
Following Lehnert et al. (2024), we consider the Maze and the Sokoban tasks and use the same dataset. Both datasets contain training examples with complete A* search trace. The A* implementation is non-deterministic where it breaks cost ties randomly and randomizes the order in which child nodes are expanded. The size of Maze varies from to . For all the Maze tasks, we randomly generate percentage of wall cells as obstacles, and randomly sample the goal and start locations. The Sokoban map is of size with randomly placed two docks, two boxes, and the worker location. We also randomly add two additional wall cells to the interior of the map. For the detailed map generation procedure and example figures, we refer the readers to Appendix 7.1.
We first demonstrate that Dualformer can be explicitly controlled to operate in either fast or slow mode in Section 4.1. It will only output the final plan in fast mode, while a reasoning trace will be generated in slow mode. In Section 5.1.1-5.1.2, we compare Dualformer with the corresponding baselines in each mode, respectively. A variety of metrics, as listed below, are used to systematically evaluate the performance, including the correctness, optimality and diversity of generated plans, length of the reasoning trace, etc. Last, we ablate our design choices in Section 5.1.5.
@@
We instantiate Dualformer using the same encoder-decoder architecture as in Lehnert et al. (2024). The encoder is an adaptation of the T5 architecture Raffel et al. (2020) with rotary embeddings, while the decoder is a GPT style architecture. We employ a model size of 15M and 46M for the Maze and the Sokoban environments, respectively. All models are trained on 100k training examples. We trained the model for iterations for the Maze environment and iterations for the Sokoban environment. We defer the details of model architectures and the other hyperparameters to Appendix 7.

### v3

# Dualformer: Controllable Fast and Slow Thinking by Learning with Randomized Reasoning Traces

@@ Abstract
In cognition theory, human thinking is governed by two systems: the fast and intuitive System 1 and the slower but more deliberative System 2. Analogously, Large Language Models (LLMs) can operate in two reasoning modes: outputting only the solutions (fast mode) or both the reasoning chain and the final solution (slow mode). We present Dualformer, a single Transformer model that seamlessly integrates both the fast and slow reasoning modes by training on randomized reasoning traces, where different parts of the traces are strategically dropped during training. At inference time, Dualformer can be easily configured to execute in either fast or slow mode, or automatically decide which mode to engage (auto mode). It outperforms baselines in both performance and computational efficiency across all three modes: (1) in slow mode, Dualformer achieves optimal rate on unseen maze tasks, surpassing the Searchformer baseline (93.3%) trained on data with complete reasoning traces, with fewer reasoning steps; (2) in fast mode, Dualformer achieves optimal rate, significantly outperforming the Solution-Only model trained on solution-only data, which has an optimal rate of only 30%; (3) in auto mode, Dualformer achieves optimal rate with fewer steps than Searchformer. Moreover, Dualformer produces more diverse reasoning traces than Searchformer. For math reasoning problems, our techniques have also achieved improved performance with LLM fine-tuning, demonstrating its generalization beyond task-specific models. We open source our code at https://github.com/facebookresearch/dualformer.
@@ 1 Introduction
Transformers Vaswani et al. (2017), the sequence modeling tool that serves as the cornerstone of foundation models in various domains including Large Language models (LLMs) Dosovitskiy (2020); Baevski et al. (2020); Radford et al. (2021); Touvron et al. (2021); Hsu et al. (2021); Touvron et al. (2023); Dubey et al. (2024), have been widely used in many works to approach reasoning and planning problems, see e.g., Zhou et al. (2022); Kojima et al. (2022); Pallagani et al. (2022); Valmeekam et al. (2023a); Chen et al. (2024); Gundawar et al. (2024); Wang & Zhou (2024). Specifically, we can categorize the reasoning modes of Transformers into fast and slow. In fast mode, a Transformer will output a final solution without any reasoning steps, whereas the intermediate steps of thinking, such as a search trace for finding a short path, will be generated along with the plan in slow mode.
The two inference modes share a lot of similarities with the two thinking systems inherent in us Wason & Evans (1974); Kahneman (2017): an automatic and unconscious System 1 and a controlled and conscious System 2. More importantly, they come with analogous pros and cons. As discussed in previous works Wei et al. (2022); Valmeekam et al. (2023b); Lehnert et al. (2024); Gandhi et al. (2024); Saha et al. (2024), Transformer models operating in fast mode have a lower computational cost and allow for a quicker response, yet they fall short in accuracy and optimality compared with slow mode models (see Figure B.1 for a concrete example). This raises an interesting question:
@@
Can we integrate both the fast and the slow modes into Transformer-based reasoning models, similar to how humans possess two distinct thinking systems, and let them complement each other?
@@
Multiple approaches have been proposed. A popular paradigm is to start with a pure System 2, and then fine-tune to make it more efficient like System 1, e.g., to distill System 2 output into System 1 Yu et al. (2024); Deng et al. (2024), or to improve the reasoning efficiency of existing system 2 by distillation Wang et al. (2023); Shridhar et al. (2023) or bootstrapping from symbolic systems (e.g., Searchformer Lehnert et al. (2024), Stream of Search Gandhi et al. (2024)). However, in these cases, further fine-tuning is needed, which is computationally expensive, and it is non-trivial to adapt the resulting system to be more like System 1 or System 2 on the fly. To address this problem, Saha et al. (2024) design an explicit meta-controller to switch between two different systems.
@@
We first apply our framework to train an encoder-decoder Transformer model to solve pathfinding problems, where the trace is generated by the A* search algorithm. We consider two domains: the Maze navigation and the Sokoban game as in Lehnert et al. (2024), where we use the same tokenization scheme. Interestingly, we have found that these problems are challenging for state-of-the-art LLMs like o1-preview and o1-mini, where the output path often breaks into the walls (See Appendix I for an example). In each reasoning mode, Dualformer outperforms the established baselines, achieving stronger results in both the solved rate and optimal rate. Moreover, Dualformer significantly enhances the diversity of the generated plans by identifying a greater variety of unique paths that reach the goal. Notably, Dualformer is also efficient even when working in slow mode, generating much shorter reasoning traces than the baseline model. Next, we apply our framework to fine-tune LLMs for answering math questions. Following Yu et al. (2023), the training examples are taken from the MATH dataset Hendrycks et al. (2021) where answers are rewritten by a Llama-3.1-70B-Instruct model to include detailed intermediate steps. Likewise, the obtained LLMs demonstrate enhanced efficacy and efficiency.
@@ 2 Related Work
Tremendous efforts have been made to enhance the capability of Transformer-based models to plan and reason over a long horizon. Two main types of approaches have been developed. The first type leverages existing LLMs. For instance, researchers taught LLMs to call external existing symbolic solvers, such as those found in Ahn et al. (2022); Besta et al. (2024); Sel et al. (2023); He-Yueya et al. (2023); Liu et al. (2023); Silver et al. (2024). Pallagani et al. (2022; 2024) investigate the fine-tuning of LLMs to use a symbolic solver, Schick et al. ; Hao et al. (2024) fine-tune LLMs to use external tools, and Hao et al. propose to fine-tune LLMs to do reasoning and planning within a world model. Among them, several works try to integrate System-1 and System-2 thinking into LLM reasoning. Weston & Sukhbaatar (2023) proposed the System 2 Attention (S2A), a more deliberate attention mechanism aimed at improving the reasoning capabilities of large language models by reducing their susceptibility to spurious correlations in the context. Yu et al. (2024) distill System 2 output into a System 1 model, which aims to compile higher-quality outputs from System 2 techniques back into LLM generations without intermediate reasoning token sequences.
@@
Large language models (LLMs) have been utilized for synthetic data generation in various domains. For instance, Wei et al. (2021); Longpre et al. (2023); Taori et al. (2023) introduced a synthetic instruction dataset by sampling from diverse templates containing natural language instructions that outline specific tasks. This approach has also been applied to the visual domain Liu et al. (2024b; a); Zhu et al. (2023); Brooks et al. (2023); Peng et al. (2023). To improve the performance of LLM to answer math questions, Yu et al. (2023) developed a method to rewrite, verify, and augment the original MATH dataset Hendrycks et al. (2021) using specialized prompts. Similar methodologies have been further explored in other studies, including those by Yuan et al. (2023); Luo et al. (2023); Lee et al. (2023); Yue et al. (2023); Tong et al. (2025).
@@ 3 Structured Trace Dropping and Randomized Training
Our work builds upon the Searchformer work of Lehnert et al. (2024). To perform planning, we train a Transformer to model a sequence of tokens that sequentially represents the planning task (prompt), the computation of A* algorithm (trace), and the optimal path (solution) derived. See Appendix A to the details of our tokenization scheme and example input for a Maze navigation task.
Searchformer has proven efficacy in addressing a variety of complex decision-making tasks. However, it still suffers from two important limitations. Firstly, the model only operates in slow mode and outputs lengthy reasoning chains, which significantly increases the inference time. While this can be reduced by bootstrapping Lehnert et al. (2024), an iterative refining technique that consists of cycles of rollouts followed by fine-tuning, such a procedure incurs significant extra demand on computational resources. Secondly, Searchformer struggles to generate diverse solutions, where identical rollouts are frequently sampled111We note that the lack of diversity is not due to the limitation of training data. For all our experiments, the search traces in the training datasets are generated by non-deterministic A* algorithm that randomly breaks cost ties.. For example, across 1000 maze problems we have tested, Searchformer’s reasoning chain contains more than 1500 tokens on average, and can only find unique feasible paths out of responses (see Section 4.1.2).
To address these challenges, we propose a training framework that utilizes randomized reasoning traces. Our approach is inspired by two lines of work. First, we have noticed that even though Searchformer is trained on complete A* search traces, it generates shorter traces that are sketching the search process. Second, research has shown that humans often rely on shortcuts and patterns when making decisions, a concept known as System 1 thinking Kahneman (2017). These observations, combined with the success of dropout technique Hinton (2012); Srivastava et al. (2014) that randomly drop units from the neural network during training, motivated us to explore the use of randomized reasoning traces in our framework, and we aim to simplify the A* search trace by exploiting its structured elements and selectively dropping certain parts for each training example.
@@
Level 1 strategy eliminates all the close clauses from a search trace.
Level 2 strategy goes further by additionally dropping all the cost tokens.
Level 3 strategy is even more aggressive. It further randomly drops of the create clauses.
Level 4 strategy drops the entire trace.
@@
Curious readers might already be wondering that whether our training framework resembles the token masking techniques used by famous LLMs including BERT Devlin (2018); Liu (2019); Song (2019); Gauthier & Levy (2019); Sinha et al. (2021); Kitouni et al. (2024). However, there are significant differences that distinguish our approach from those masking techniques. First, standard masking techniques usually mask the input tokens of a sequence uniformly in random. In contrast, our dropping strategies only apply to the search trace. Second, while masked LLMs generally employ bidirectional attention layers and predict the masked tokens, Dualformer uses causal attention layers, and our training objective solely focuses on the next token prediction with the overall goal of improving its reasoning and planning capability. Computationally, our training procedure is also more efficient. Dropping tokens shortens the input sequence and saves computation. For example, it takes 30 hours to train Dualformer for the maze task on 8 Tesla V100 32GB GPUs, while it requires 36 hours if we use the full reasoning trace. We defer the training details to Section 4 and Appendix C.
@@ 3.1 Controllable Generation
One appealing property of Dualformer is that it can be easily prompted to operate in either fast or slow generation mode at inference time. The control mechanism is extremely simple: we append bos and a control token to the standard prompt (which includes environment and task description), where the control token is either plan or create. If we use plan, Dualformer will operate in fast mode and directly output the plan, bypassing the reasoning steps. On the other hand, if we inject create after bos, Dualformer will work in slow mode and generate both the reasoning trace and the final plan. See Appendix D for concrete examples. If we only use the standard prompt, Dualformer will mimic the dual process of human decision making—depending on the situation, it generates either types of responses, which correspond to System 1 and System 2 reasoning.
@@ 4 Experiments
In slow and auto model, does Dualformer lead to faster reasoning, i.e., output a shorter trace?
@@
We answer questions 1 and 2 in Section 4.1, where we train Transformers to solve Maze navigation tasks and Sokoban games, similar to Searchformer Lehnert et al. (2024). To answer questions 3, we fine-tune LLama-3.1-8B and Mistral-7B models to solve math problems in Section 4.2.
@@ 4.1 Navigation Tasks: Maze & Sokoban
Following Lehnert et al. (2024), we consider the Maze and the Sokoban tasks and use the same dataset. Both datasets contain training examples with complete A* search trace. The A* implementation is non-deterministic where it breaks cost ties randomly and randomizes the order in which child nodes are expanded. The size of Maze varies from to . For all the Maze tasks, we randomly generate percentage of wall cells as obstacles, and randomly sample the goal and start locations. The Sokoban map is of size with randomly placed two docks, two boxes, and the worker location. We also randomly add two additional wall cells to the interior of the map. For the detailed map generation procedure and example figures, we refer the readers to Appendix C.1.
We first demonstrate that Dualformer can be explicitly controlled to operate in either fast or slow mode in Section 3.1. It will only output the final plan in fast mode, while a reasoning trace will be generated in slow mode. In Section 4.1.1-4.1.2, we compare Dualformer with the corresponding baselines in each mode, respectively. A variety of metrics, as listed below, are used to systematically evaluate the performance, including the correctness, optimality and diversity of generated plans, length of the reasoning trace, etc. Last, we ablate our design choices in Section 4.1.5.
@@
We instantiate Dualformer using the same encoder-decoder architecture as in Lehnert et al. (2024). The encoder is an adaptation of the T5 architecture Raffel et al. (2020) with rotary embeddings, while the decoder is a GPT style architecture. We employ a model size of 15M and 46M for the Maze and the Sokoban environments, respectively. All models are trained on 100k training examples. We trained the model for iterations for the Maze environment and iterations for the Sokoban environment. We defer the details of model architectures and the other hyperparameters to Appendix C.